calculate the vapour pressure of a solvent above a solution containing 83.99g of solute of molecular mass 195 .62per 1000g of water 20°c is 17.53 mmhg

1
Expert's answer
2022-02-03T15:15:01-0500

pH2O=pH2O**xH2O (Raoult's law)

nH2O=1200/18 = 66.667 mol

nx = 83.99/195.62 = 0.429

xH2O = 66.667/(66.667+0.429) = 0.99361

pH2O=17.53×0.99361 = 17.42 mmhg